OpenStudy (anonymous):

Calculate the pH of a solution prepared by mixing 15.0 mL of 0.10 M NaOH and 30.0 mL of a 0.10 M benzoic acid solution. (Benzoic acid is monprotic; its dissociation constant is 6.5 x 10^-5.)

OpenStudy (anonymous):

Balance Equation C65COOH(aq) +OH(aq) -> C6H5COO+ Na(aq) +H20(l) 1.) Get units of Molar Concentration Moles NaOH= ( 0.015L)(0.10) = 0.0015M NaOH Moles C6H5COOH= ( 0.030L)( 0.10)= 0.030M C6H5COOH Benzoic Acid 2.) Find total concentration of C6H5COOH, this will determine if this solution is a buffer C6H5COOH= 0.0015mol/ 0.0450L = 0.0333M C6H5COO = 0.0015mol/ 0.00450L= 0.0333M since they are equal concentrations, this means this is a Buffer 3.) PH=Pka+log Benzoic acid/ Benzoic Ion so you can just do this pH= -log( 6.5*10^-5) pH= 4.19

OpenStudy (anonymous):

Let me know if you understand

OpenStudy (anonymous):

@n648c788 i understand but doesn't (0.030L)(0.10)= 0.0030M NOT 0.030M ?

OpenStudy (anonymous):

yes I forgot an extra o sorry, but that is how you do it

OpenStudy (anonymous):

ok, so the answer would not be 4.19?

OpenStudy (anonymous):

It will be 4.19, because the extra work is to prove it is a buffer, so just take the -log of the 6.5*10^-5

OpenStudy (anonymous):

If I calculated part 2 wrong then we would have a different answer, but I did part 2 correctly it showed that this reaction is a buffer, so we were able to take the -log of the ka

OpenStudy (anonymous):

Just notfity I, if you still need clarification the answer is 4.19

OpenStudy (anonymous):

pH= pka +log conj acid/base pka=-log pH=-log(6.5*10^-5)+log(0.0333M/0.0333M) pH=4.19+0 pH=4.19